Can you use the normal distribution to approximate the binomial distribution. Give reason?
Yes, and the justification comes from the Central Limit Theorem.
How do you find the semi-interquartile range with only the mean and standard deviation?
In general, you cannot.
If the distribution can be assumed to be Gaussian [Normal] then you could use z-scores.
What is probability explain conditional probability?
Probability is the chance of an event occurring. For example when flipping a coin you have a 50% chance that it will land on heads and a 50% chance that it will land on tails since there are only two possibilities.
Conditional probability refers to when one event is dependent on another event occurring. It can also be written as the probability of an event B occurring after event A has already occurred. The notation for conditional probability is P(B|A). (Note: this does not mean B divided by A but probability of B after A)
When two events are dependent, the probability of them both occurring is:
P(A and B)=P(A)P(B|A)
So for example: 53% of residents have home owners insurance. Of them, 27% has auto-insurance. If a resident is selected at random, what is the probability they with have both insurances?
Let H stand for home owners insurance = 53% or 0.53
Let A stand for auto insurance = 27% or 0.27
P(H and P)=P(H)P(A/H)
=(0.53)(0.27)
= 0.1431
So the probability of residents have both home owners and auto insurance is 0.1431 or 14.31%
The answer is 11C2 * 8C3 = 55*56 = 3080 ways.
How can you use odd to find probability?
Odds of A to B in favour of an event states that for every A times an event occurs, the event does not occur B times. So, out of (A+B) trials, A are favourable to the event. that is, the probability of A is A/(A+B).
How many simple random samples of size 5 can be selected from a population of size 35?
There are 324,632 possible samples.
The chance that a given event will occur usually expressed between the number 0 will not and 1 the event will occur
What is the probability of pulling a 5 or a heart from a deck of cards?
Well, if you include the jokers, you have 52 cards total. There are four suits, so that's four "5" cards in the deck, which gives you a 4 in 52 (or 2 in 21, simplified) probability. Each suit has 13 cards, so you have a 13 in 52 chance of pulling a card of any particular suit (13 is a prime number, and can't be simplified).
What does data distribution mean?
It is the set of values that a variable can take together with the probability or frequency distribution for those values.
What is the probabilty of drawing a face card that has a diamond?
3/52
Assuming you're randomly drawing from a usual deck 52 unique cards. Jack Queen King are the face cards.
How many distinct arrangements can be made with the letters in the word banana?
There are 6!/(3!*2!) = 60 arrangements.
What is the probability of getting a heart out of a pack of cards?
If the pick is completely random, the deck is a standard deck and there are no jokers or any other cards other than the standard 52, the probability is 1/4
It is 1/13*1/51 = 1/663 = 0.001508 approx.
25.00
There are 11880 ways.
It is: 12 out of 50 or 6 out of 25
What is the probability to take a Jack in 52 cards?
1:13
or 4:52
as four suits and one jack in each suit
If there is a correlation between event a and Event B what statements is true?
event B has something in common with event A
Answer: 1/10 for no tickets, 6/10 for two tickets, 3/10 tickets for one ticket.
What are two possible outcomes to prove an experiments hypothesis?
To prove the hypothesis.
To disprove the hypothesis.